# Repetitive Contrastive Learning Enhances Mamba’s Selectivity in Time Series Prediction

### v1

@@ 4.1.2 Temporal Baseline
    Transformer: (Vaswani et al., 2023) The Transformer model, introduced by Vaswani et al. in 2017, revolutionized sequence modeling by using self-attention mechanisms. Its architecture allows for efficient parallelization and effectively captures long-range dependencies, making it highly suitable for various tasks such as natural language processing and time series forecasting.
@@
    iTransformer:(Liu et al., 2024) iTransformer is a restructured Transformer tailored for multivariate time series forecasting. Instead of embedding simultaneous time steps, it encodes each variate’s full time series as a token, enabling better capture of global patterns and cross-variate correlations. This design aligns attention with the intrinsic structure of time series and achieves strong performance across forecasting benchmarks.
@@
    TimeMixer: (Wang et al., 2024) TimeMixer is a novel approach designed for time series modeling, leveraging the power of mixing operations to combine temporal features. By focusing on capturing intricate temporal dependencies and interactions, TimeMixer provides robust performance in both short-term and long-term forecasting tasks.
@@
    N-Beats: (Oreshkin et al., 2020) N-Beats is a deep learning architecture designed for univariate time series forecasting. It employs a Doubly Residual Stacking mechanism, utilizing both backward and forward residual links to enhance signal propagation through a deep stack of fully connected layers. The model is highly flexible, requiring no time-series-specific components, and demonstrates state-of-the-art performance across diverse datasets, including M3, M4, and TOURISM. Additionally, N-Beats incorporates Ensembling techniques during training, further improving its robustness and accuracy.
@@
    N-HiTS: (Challu et al., 2022) N-HiTS builds upon N-Beats by introducing a Neural Basis Approximation Theorem to enhance theoretical guarantees in forecasting. The model significantly improves long-horizon predictions through Multi-Rate Signal Sampling, allowing it to focus on different frequency components dynamically. Additionally, N-HiTS employs a Hierarchical Interpolation mechanism, which enables efficient decomposition and synthesis of forecasted signals, reducing volatility and computational complexity.
@@
    CrossFormer: (Wang et al., 2021) CrossFormer introduces a cross-attention mechanism specifically tailored for time series data. It excels in integrating multiple time series inputs, enabling the model to learn complex relationships across different temporal sequences, thus improving forecasting accuracy and adaptability to diverse datasets.
@@
    PatchTST: (Nie et al., 2023) PatchTST is a model that applies the concept of patch-based processing from computer vision to time series data. By segmenting time series into patches and processing them independently, PatchTST enhances the model’s ability to capture local temporal patterns, improving efficiency and scalability for large datasets.
@@
    TimesNet: (Wu et al., 2023) TimesNet is an advanced time series network that leverages a hierarchical structure to model temporal dependencies at multiple scales. This architecture allows TimesNet to adaptively focus on different temporal resolutions, providing superior performance in multiscale time series forecasting.
@@
    FEDFormer: (Zhou et al., 2022) FEDFormer incorporates federated learning principles into the Transformer framework, allowing for decentralized time series modeling. This model is particularly effective in scenarios where data privacy is crucial, as it can learn from distributed data sources without centralizing the datasets.
@@
    Informer: (Zhou et al., 2021) Informer is designed to efficiently handle long sequences in time series forecasting. It introduces a ProbSparse self-attention mechanism that reduces computational complexity and memory usage, making it ideal for real-time applications and large-scale datasets. Informer achieves state-of-the-art results by focusing on significant temporal patterns while filtering out noise.
@@ 4.1.3 Temporal Pre-training Baseline
    SoftCLT: (Lee et al., 2024) SoftCLT is a cutting-edge model designed for contextual sequence learning. By incorporating soft clustering techniques, SoftCLT dynamically groups similar temporal patterns, enhancing the model’s ability to generalize across varied contexts. This approach ensures superior performance in complex classification tasks, offering robust adaptability to fluctuating sequences while maintaining high interpretability.
@@
    InfoTS: (Luo et al., 2023) InfoTS leverages information-theoretic principles to optimize time series modeling. By prioritizing the retention of informative features and minimizing redundancy, InfoTS significantly enhances predictive accuracy. This model excels in both supervised and unsupervised learning scenarios, making it versatile for diverse applications such as anomaly detection and trend analysis.
@@ 4.1.6 Model Settings
The parameter settings for the Mamba block during pre-training are as follows: The model dimension ($d_{model}$) is set to values [16, 32, 64], and the state dimension ($d_{state}$) is set to [16, 64, 128]. The convolution dimension ($d_{conv}$) is fixed at 4, and $pad\_vocab\_size\_multiple$ is set to 8 to ensure consistent padding sizes. The expansion factor ($expand$) is configured to 2, with $conv\_bias$ enabled (set to True) and $bias$ disabled (set to False). The repeat time, denoted as $n_{t}$, is set to 3, while noise variance is varied between [0.001, 0.01]. During the inference phase, the Mamba Selective State Space Model (SSM) parameters are aligned with the corresponding pre-trained block parameters to maintain consistency and leverage learned patterns effectively.
@@ 4.1.7 Training Settings
The experiment was conducted on a server equipped with four NVIDIA GeForce RTX 3090 GPUs and an AMD EPYC 7282 16-Core Processor. During the pre-training phase, the number of layers ($n\_layer$) is set to 1, the number of epochs ($epoch$) is 100, the learning rate ($lr$) is configured to 1e-4, and the regularization coefficient is also set to 1e-4. In the inference stage, the maximum number of training epochs remains at 100, while $n\_layer$ is increased to 4. The Mean Absolute Error (MAE) serves as the loss function, and model selection is based on the lowest validation set loss. The parameter $frozentype$ is chosen as needed from the options [None, FrozenA], and the number of layers used for parameter replacement is selected from [25%, 50%, 75%, 100%], according to the specific experimental requirements. For the prediction length, we selected four different lengths: [96, 192, 336, 720] and conducted a series of experiments. However, all results tables presented in our paper, unless otherwise specified, use a prediction length of 96. This length was chosen because it effectively illustrates the corresponding conclusions and provides a clear basis for our findings.
@@ 4.3 Comparison with Pre-training Methods
We conducted a series of experiments on the latest pre-training methods in the time series domain (Luo et al., 2023; Lee et al., 2024). The results, presented in Table 2, were derived from models trained using official code on multivariate forecasting tasks. Two important aspects warrant attention. First, both methods are designed to enhance the representation learning of time series features through contrastive pre-training, heavily relying on the capabilities of feature extraction modules. Specifically, their experiments utilized the TSEncoder from woTS2Vec(Yue et al., 2022) or TC from CATCC(Eldele et al., 2023) as feature extractors. These models are structurally distinct from mamba-based models, leading to a decline in performance when feature extraction is adapted to mamba models. Second, these approaches primarily benefit classification tasks due to their ability to accurately and effectively represent time series nodes, which aids classification but demonstrates limited improvements in forecasting tasks especially in multivariate tasks. Consequently, during their prediction stages, they use feature vectors from pre-training train a linear model to predict future values instead of leveraging pre-trained modules to construct new models. In contrast, our pre-training approach guides mamba blocks to learn sampling rules inherent in natural time sequences and identify meaningful historical information. This aligns with the requirements of forecasting tasks, allowing us to directly leverage parameters in forecasting models for superior results.
Model TimeMachine* Bi-Mamba* Mamba* iMamba* InfoTS(TS2Vec) SoftCLT(TS2Vec) SoftCLT(Mamba) InfoTS(Mamba)
Metric MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE
ETTh1 96 0.387 0.379 0.389 0.379 0.575 0.657 0.499 0.493 0.623 0.736 0.616 0.704 0.696 0.891 0.816 1.147
192 0.420 0.440 0.421 0.425 0.602 0.713 0.508 0.532 0.690 0.857 0.670 0.810 0.737 0.959 0.835 1.186
336 0.442 0.482 0.456 0.481 0.608 0.715 0.513 0.550 0.769 1.024 0.740 0.950 0.640 1.064 0.861 1.231
ETTh2 96 0.330 0.282 0.347 0.300 1.228 2.124 0.693 0.908 0.754 0.936 0.799 1.015 0.997 1.542 0.897 1.219
192 0.382 0.355 0.394 0.373 1.237 2.164 1.023 1.821 1.112 2.022 1.251 2.559 1.343 2.820 1.251 2.506
336 0.420 0.412 0.429 0.434 1.234 2.153 1.073 2.042 1.264 2.482 1.312 2.639 1.402 2.952 1.327 2.733
ETTm1 96 0.346 0.318 0.358 0.332 0.492 0.528 0.432 0.400 0.540 0.602 0.534 0.581 0.623 0.808 0.741 0.985
192 0.377 0.375 0.384 0.369 0.513 0.587 0.450 0.439 0.575 0.649 0.569 0.635 0.654 0.849 0.756 1.014
336 0.387 0.396 0.407 0.404 0.817 1.457 0.491 0.509 0.622 0.729 0.610 0.697 0.681 0.885 0.770 1.040
ETTm2 96 0.251 0.173 0.271 0.186 0.576 0.601 0.416 0.367 0.452 0.377 0.460 0.400 0.491 0.437 0.782 0.969
192 0.293 0.238 0.313 0.254 0.667 0.847 0.497 0.495 0.560 0.542 0.580 0.587 0.591 0.590 0.857 1.152
336 0.333 0.299 0.364 0.316 0.705 0.922 0.793 1.032 0.713 0.846 0.730 0.885 0.730 0.855 0.969 1.461
Electricity 96 0.259 0.183 0.261 0.182 0.423 0.393 0.260 0.183 0.290 0.380 0.401 0.326 0.553 0.571 0.531 0.524
192 0.246 0.152 0.270 0.188 0.430 0.405 0.280 0.205 0.293 0.383 0.403 0.327 0.555 0.573 0.532 0.524
336 0.261 0.169 0.283 0.200 0.435 0.411 0.298 0.222 0.311 0.396 0.416 0.344 0.565 0.581 0.540 0.538
@@ 4.6 Comparison with Temporal Model
We compared our approach with existing state-of-the-art time series prediction models. We set all input lengths to 96 and conducted experiments across multiple prediction horizons $T=\{96,192,336,720\}$. The average results across the four prediction horizons are presented in Table 3, while the detailed results for each individual prediction horizon are provided in Table 10 of B.1. TimeMachine* and Bi-Mamba* refer to the TimeMachine and Bi-Mamba models initialized with parameters obtained using RCL. Our method achieves optimal results across various datasets and prediction horizons. For datasets with fewer data channels, our approach consistently achieves the best Mean Absolute Error (MAE) results across all prediction horizons, and Mean Squared Error (MSE) results are generally among the top two. For datasets with more channels, such as traffic and electricity, our method shows more significant improvements for longer prediction targets. This indicates enhanced stability in long-sequence predictions, attributed to the parameters obtained through RCL, which enable the Mamba block to have stronger selectivity for time series data.

### v2

@@ 4.1.2 Temporal Baseline
    Transformer: (vaswani2023attentionneed) The Transformer model, introduced by Vaswani et al. in 2017, revolutionized sequence modeling by using self-attention mechanisms. Its architecture allows for efficient parallelization and effectively captures long-range dependencies, making it highly suitable for various tasks such as natural language processing and time series forecasting.
@@
    iTransformer:(liu2024itransformerinvertedtransformerseffective) iTransformer is a restructured Transformer tailored for multivariate time series forecasting. Instead of embedding simultaneous time steps, it encodes each variate’s full time series as a token, enabling better capture of global patterns and cross-variate correlations. This design aligns attention with the intrinsic structure of time series and achieves strong performance across forecasting benchmarks.
@@
    TimeMixer: (wang2024timemixerdecomposablemultiscalemixing) TimeMixer is a novel approach designed for time series modeling, leveraging the power of mixing operations to combine temporal features. By focusing on capturing intricate temporal dependencies and interactions, TimeMixer provides robust performance in both short-term and long-term forecasting tasks.
@@
    N-Beats: (oreshkin2020nbeatsneuralbasisexpansion) N-Beats is a deep learning architecture designed for univariate time series forecasting. It employs a Doubly Residual Stacking mechanism, utilizing both backward and forward residual links to enhance signal propagation through a deep stack of fully connected layers. The model is highly flexible, requiring no time-series-specific components, and demonstrates state-of-the-art performance across diverse datasets, including M3, M4, and TOURISM. Additionally, N-Beats incorporates Ensembling techniques during training, further improving its robustness and accuracy.
@@
    N-HiTS: (challu2022nhitsneuralhierarchicalinterpolation) N-HiTS builds upon N-Beats by introducing a Neural Basis Approximation Theorem to enhance theoretical guarantees in forecasting. The model significantly improves long-horizon predictions through Multi-Rate Signal Sampling, allowing it to focus on different frequency components dynamically. Additionally, N-HiTS employs a Hierarchical Interpolation mechanism, which enables efficient decomposition and synthesis of forecasted signals, reducing volatility and computational complexity.
@@
    CrossFormer: (wang2021crossformerversatilevisiontransformer) CrossFormer introduces a cross-attention mechanism specifically tailored for time series data. It excels in integrating multiple time series inputs, enabling the model to learn complex relationships across different temporal sequences, thus improving forecasting accuracy and adaptability to diverse datasets.
@@
    PatchTST: (nie2023timeseriesworth64) PatchTST is a model that applies the concept of patch-based processing from computer vision to time series data. By segmenting time series into patches and processing them independently, PatchTST enhances the model’s ability to capture local temporal patterns, improving efficiency and scalability for large datasets.
@@
    TimesNet: (wu2023timesnettemporal2dvariationmodeling) TimesNet is an advanced time series network that leverages a hierarchical structure to model temporal dependencies at multiple scales. This architecture allows TimesNet to adaptively focus on different temporal resolutions, providing superior performance in multiscale time series forecasting.
@@
    FEDFormer: (zhou2022fedformerfrequencyenhanceddecomposed) FEDFormer incorporates federated learning principles into the Transformer framework, allowing for decentralized time series modeling. This model is particularly effective in scenarios where data privacy is crucial, as it can learn from distributed data sources without centralizing the datasets.
@@
    Informer: (zhou2021informerefficienttransformerlong) Informer is designed to efficiently handle long sequences in time series forecasting. It introduces a ProbSparse self-attention mechanism that reduces computational complexity and memory usage, making it ideal for real-time applications and large-scale datasets. Informer achieves state-of-the-art results by focusing on significant temporal patterns while filtering out noise.
@@ 4.1.3 Temporal Pre-training Baseline
    SoftCLT: (softcls) SoftCLT is a cutting-edge model designed for contextual sequence learning. By incorporating soft clustering techniques, SoftCLT dynamically groups similar temporal patterns, enhancing the model’s ability to generalize across varied contexts. This approach ensures superior performance in complex classification tasks, offering robust adaptability to fluctuating sequences while maintaining high interpretability.
@@
    InfoTS: (infots) InfoTS leverages information-theoretic principles to optimize time series modeling. By prioritizing the retention of informative features and minimizing redundancy, InfoTS significantly enhances predictive accuracy. This model excels in both supervised and unsupervised learning scenarios, making it versatile for diverse applications such as anomaly detection and trend analysis.
@@ 4.1.6 Model Settings
The parameter settings for the Mamba block during pre-training are as follows: The model dimension ($d_{model}$) is set to values [16, 32, 64], and the state dimension ($d_{state}$) is set to [16, 64, 128]. The convolution dimension ($d_{conv}$) is fixed at 4, and pad_vocab_size_multiplepad\_vocab\_size\_multiple is set to 8 to ensure consistent padding sizes. The expansion factor (expandexpand) is configured to 2, with conv_biasconv\_bias enabled (set to True) and biasbias disabled (set to False). The repeat time, denoted as $n_{t}$, is set to 3, while noise variance is varied between [0.001, 0.01]. During the inference phase, the Mamba Selective State Space Model (SSM) parameters are aligned with the corresponding pre-trained block parameters to maintain consistency and leverage learned patterns effectively. When comparing with other models and pre-training methods, we uniformly set $d_{model}=64$ and $d_{state}=16$. For the parameters of RCL, we choose to repeat each time step $n_{t}=3$ times, and the noise distribution is set to $\sigma_{a}=1e-3,\sigma_{b}=1e-2$.
@@ 4.1.7 Training Settings
The experiment was conducted on a server equipped with four NVIDIA GeForce RTX 3090 GPUs and an AMD EPYC 7282 16-Core Processor. We conducted each experiment at least three times and provided the standard deviation of the results in D.4. During the pre-training phase, the number of layers (n_layern\_layer) is set to 1, the number of epochs (epochepoch) is 100, the learning rate ($lr$) is configured to 1e-4, and the regularization coefficient is also set to 1e-4. In the inference stage, the maximum number of training epochs remains at 100, while n_layern\_layer is increased to 4. The Mean Absolute Error (MAE) serves as the loss function, and model selection is based on the lowest validation set loss. The parameter frozentypefrozentype is chosen as needed from the options [None, FrozenA], and the number of layers used for parameter replacement is selected from [25%, 50%, 75%, 100%], according to the specific experimental requirements. For the prediction length, we selected four different lengths: [96, 192, 336, 720] and conducted a series of experiments. However, all results tables presented in our paper, unless otherwise specified, use a prediction length of 96. This length was chosen because it effectively illustrates the corresponding conclusions and provides a clear basis for our findings.
@@ 4.3 Comparison with Pre-training Methods
To systematically evaluate the latest pretraining methods in the time series domain, we conducted experiments on TS2Vec (yue2022ts2vecuniversalrepresentationtime) and InfoTS (infots), and SoftCLT(softcls), with results summarized in Table 2. These experiments utilized official code implementations and focused on multivariate forecasting tasks. Below, we compare these contrastive methods based on two key dimensions: design goals, and experiments performance, contrasting these with our proposed pretraining approach tailored for Mamba-based models.
TS2Vec(yue2022ts2vecuniversalrepresentationtime), InfoTS(infots), and SoftCLTsoftcls aim to enhance time series representation learning via contrastive pretraining, focusing on robust feature extraction for classification tasks. TS2Vec(yue2022ts2vecuniversalrepresentationtime) employs hierarchical, dual-objective contrastive learning for generalizable embeddings, InfoTS(infots) maximizes mutual information for discriminative representations, and SoftCLT(softcls) leverages soft contrastive objectives to improve feature robustness. However, these methods, optimized for classification, rely on linear models for forecasting, limiting their efficacy in multivariate settings. In contrast, our RCL pretraining for Mamba-based models targets forecasting by learning intrinsic temporal patterns, enabling direct integration of pretrained parameters and achieving superior performance in forecasting tasks.
Our RCL pretraining approach, when applied to Mamba-based models, achieves a performance improvement of approximately 20% in time series forecasting tasks compared to Mamba models pretrained with InfoTS, TS2Vec, or SoftCLT, as demonstrated by experimental results in Table 2.Unlike existing methods such as TS2Vec (yue2022ts2vecuniversalrepresentationtime) and InfoTS (infots), which exhibit degraded performance when adapted to Mamba architectures due to their reliance on structurally incompatible feature extractors like TSEncoder, our method innovatively aligns pretraining with the linear processing strengths of Mamba models. By focusing on learning intrinsic temporal sampling rules and historical patterns through tailored contrastive learning, our approach ensures robust parameter initialization that directly enhances forecasting accuracy. This novel design not only outperforms alternative pretraining strategies but also establishes a new benchmark for scalable and effective time series forecasting, particularly in multivariate settings.
Model TimeMachine* Bi-Mamba* Mamba* iMamba* TS2Vec(TSEncoder) InfoTS(TSEncoder) SoftCLS(TSEncoder) TS2Vec(Mamba) InfoTS(Mamba) SoftCLS(Mamba)
Metric MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE
ETTh1 96 0.387 0.379 0.389 0.379 0.575 0.657 0.499 0.493 0.606 0.691 0.623 0.736 0.616 0.704 0.793 1.089 0.816 1.147 0.696 0.891
192 0.420 0.440 0.421 0.425 0.602 0.713 0.508 0.532 0.663 0.798 0.690 0.857 0.670 0.810 0.825 1.149 0.835 1.186 0.737 0.959
336 0.442 0.482 0.456 0.481 0.608 0.715 0.513 0.550 0.739 0.945 0.769 1.024 0.740 0.950 0.854 1.208 0.861 1.231 0.640 1.064
avg 0.416 0.434 0.422 0.428 0.595 0.695 0.506 0.525 0.669 0.811 0.694 0.872 0.676 0.821 0.824 1.149 0.838 1.188 0.691 0.971
ETTh2 96 0.330 0.282 0.347 0.300 1.228 2.124 0.693 0.908 0.806 1.031 0.754 0.936 0.799 1.015 1.062 1.693 0.897 1.219 0.997 1.542
192 0.382 0.355 0.394 0.373 1.237 2.164 1.023 1.821 1.174 2.118 1.112 2.022 1.251 2.559 1.294 2.536 1.251 2.506 1.343 2.820
336 0.420 0.412 0.429 0.434 1.234 2.153 1.073 2.042 1.308 2.618 1.264 2.482 1.312 2.639 1.375 2.763 1.327 2.733 1.402 2.952
avg 0.377 0.350 0.390 0.369 1.233 2.147 0.929 1.590 1.096 1.922 1.043 1.813 1.121 2.071 1.244 2.331 1.158 2.152 1.248 2.438
ETTm1 96 0.346 0.318 0.358 0.332 0.492 0.528 0.432 0.400 0.530 0.572 0.540 0.602 0.534 0.581 0.774 1.031 0.741 0.985 0.623 0.808
192 0.377 0.375 0.384 0.369 0.513 0.587 0.450 0.439 0.562 0.624 0.575 0.649 0.569 0.635 0.784 1.051 0.756 1.014 0.654 0.849
336 0.387 0.396 0.407 0.404 0.817 1.457 0.491 0.509 0.603 0.686 0.622 0.729 0.610 0.697 0.802 1.088 0.770 1.040 0.681 0.885
avg 0.370 0.363 0.383 0.368 0.607 0.857 0.458 0.449 0.565 0.627 0.579 0.660 0.571 0.638 0.787 1.056 0.755 1.013 0.653 0.847
ETTm2 96 0.251 0.173 0.271 0.186 0.576 0.601 0.416 0.367 0.453 0.401 0.452 0.377 0.460 0.400 0.855 1.113 0.782 0.969 0.491 0.437
192 0.293 0.238 0.313 0.254 0.667 0.847 0.497 0.495 0.576 0.579 0.560 0.542 0.580 0.587 0.923 1.282 0.857 1.152 0.591 0.590
336 0.333 0.299 0.364 0.316 0.705 0.922 0.793 1.032 0.717 0.862 0.713 0.846 0.730 0.885 1.008 1.540 0.969 1.461 0.730 0.855
avg 0.292 0.237 0.316 0.252 0.650 0.790 0.569 0.631 0.582 0.614 0.575 0.588 0.590 0.624 0.929 1.311 0.869 1.194 0.604 0.627
Electricity 96 0.259 0.183 0.261 0.182 0.423 0.393 0.260 0.183 0.403 0.329 0.290 0.380 0.401 0.326 0.635 0.696 0.531 0.524 0.553 0.571
192 0.246 0.152 0.270 0.188 0.430 0.405 0.280 0.205 0.406 0.330 0.293 0.383 0.403 0.327 0.636 0.698 0.532 0.524 0.555 0.573
336 0.261 0.169 0.283 0.200 0.435 0.411 0.298 0.222 0.419 0.348 0.311 0.396 0.416 0.344 0.639 0.702 0.540 0.538 0.565 0.581
avg 0.255 0.168 0.271 0.190 0.429 0.403 0.279 0.203 0.409 0.335 0.298 0.386 0.407 0.332 0.637 0.699 0.534 0.529 0.558 0.575
@@ 4.6 Comparison with Temporal Model
We compared our approach with existing state-of-the-art time series prediction models. We set all input lengths to 96 and conducted experiments across multiple prediction horizons $T=\{96,192,336,720\}$. The average results across the four prediction horizons are presented in Table 3, while the detailed results for each individual prediction horizon are provided in Table 12 of D.1. TimeMachine* and Bi-Mamba* refer to the TimeMachine and Bi-Mamba models initialized with parameters obtained using RCL. Our method achieves optimal results across various datasets and prediction horizons. For datasets with fewer data channels, our approach consistently achieves the best Mean Absolute Error (MAE) results across all prediction horizons, and Mean Squared Error (MSE) results are generally among the top two. For datasets with more channels, such as traffic and electricity, our method shows more significant improvements for longer prediction targets. This indicates enhanced stability in long-sequence predictions, attributed to the parameters obtained through RCL, which enable the Mamba block to have stronger selectivity for time series data.